OpenStudy (anonymous):

Find the exact area of the minor segment of a circle if radius is 10cm and the angle subtended is 5π/6

OpenStudy (amistre64):

t r^2/2

OpenStudy (amistre64):

(5pi/6) 10(5) = 250pi/6

OpenStudy (amistre64):

125pi/3 ?

OpenStudy (anonymous):

following so far, its converting the Sin(5π/6) into an exact ratio that gets me

OpenStudy (amistre64):

so about 125 cm^2

OpenStudy (amistre64):

dont need to convert sin; it just theta/2 r^2

OpenStudy (amistre64):

the total area of a circle is 2pi/2 r^2 right?

OpenStudy (anonymous):

no this is a minor segment :/ need to use the A=(1/2)r^2(@-Sin@) (where @ = theta and in radians)

OpenStudy (amistre64):

define minor segment....is that new terminology?

OpenStudy (amistre64):

aint it just the pie shaped piece?

OpenStudy (amistre64):

ohhhh minus the triangle lol

OpenStudy (amistre64):

the triangle is side side/2 sin(t)

OpenStudy (anonymous):

Minor means small. :)

OpenStudy (anonymous):

yes but i need to use the above formula to figure out and converting the Sin part to exact ratios confuses the hell out of me

OpenStudy (amistre64):

the sin(5pi/6) = 1/2

OpenStudy (anonymous):

really... how? :/

OpenStudy (amistre64):

you take a 30 degree angle and look across it at the 1 and to the hyp of 2

OpenStudy (amistre64):

basic trig angle

OpenStudy (anonymous):

but doesn't 5pi/6 = 150 degrees? :Z

OpenStudy (amistre64):

yes; which has a reference angle of 30 degrees

OpenStudy (amistre64):

150 and 30 have the same sin

OpenStudy (anonymous):

ah yes, I get you now! cheers for the help :)

OpenStudy (amistre64):